OpenStudy (anonymous):

How would I rewrite this in vertex form and identify the vertex? y = x^2 + 4x - 6

OpenStudy (anonymous):

first divide the coefficient of x (6) by 2 which = 2and write: (x + 2)^2

OpenStudy (anonymous):

when this is expanded the last term = 4 but you want it to be -6 so u subtract 10 and the vertex form becomes (x + 2)^2 - 10

OpenStudy (anonymous):

the minimum value of the function is when (x + 2)^2 = 0 so it is -10 the x-coordinate of the minimum value is x = -2 so the minimum is at (-2,-10)

OpenStudy (amistre64):

y = x^2 + 4x - 6 y = (x^2 + 4x) - 6 ; complete the square y = (x^2 + 4x +4) -4 - 6 ; compact it and combine like terms y = (x+2)^2 -10 completeing the square would take some geometry to help understand what is going on :)

OpenStudy (anonymous):

sorry - in my first post, coefficient of x = 4 = to give (x+2)^2

OpenStudy (amistre64):

like this :)